Differentiate y= In ( 2cot²x )


1
Expert's answer
2022-03-19T02:36:59-0400

"=\frac{d}{dx}[ln(2cot^2(x))]=\frac{1}{2cot^2(x)}*\frac{d}{dx}[2cot^2(x)]"

"=\frac{2*\frac{d}{dx}[cot^2(x)]}{2cot^2(x)}"

"=\frac{2cot(x)*\frac{d}{dx}cot(x)}{cot^2(x)}"

"=\frac{2(-csc^2(x)}{cot(x)}"

"=-\frac{2csc^2}{cot(x)}"